# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In a recent paper, Li and Ju argued that generally, models of Coalition Logic, concurrent game models, have three too strong assumptions: seriality, independence of agents, and determinism. They presented a Minimal Coalition Logic based on general concurrent game models, which do not have the three assumptions. However, when constructing coalition logics about strategic reasoning in special kinds of situations, we may want to keep some of the assumptions. Thus, studying coalition logics with some of these assumptions makes good sense. In this paper, we show the completeness of these coalition logics by a uniform approach.
@@ 1 Introduction
1.1 Coalition Logic
@@
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{A}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{A}$ ensures $\phi$. Its models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; at every state, every coalition has some available joint actions; every joint action of a coalition has some possible outcome states. The formula $\langle[\mathrm{A}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{A}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
1.2 Minimal Coalition Logic
@@
Li and Ju [5] argued that generally, concurrent game models have three too strong assumptions. The first one is seriality: coalitions always have available joint actions. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: joint actions of the grand coalition always have a unique outcome. Based on general concurrent game models, which do not have the three assumptions, Li and Ju [5] presented a Minimal Coalition Logic $\mathsf{MCL}$.
1.3 It makes good sense to study the logics between Minimal Coalition Logic and Coalition Logic
@@
Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in special kinds of situations. First, consider seriality. When assuming that agents can always perform the special action doing nothing, seriality holds. Second, consider the independence of agents. In many situations, agents’ actions do not interfere with each other, and the independence of agents holds. Consider the game of rock-paper-scissors and the Prisoner’s Dilemma as examples. As mentioned by Sergot [10], when we just consider some special kind of actions, such as attempting to do something, the independence of agents does hold. Third, consider determinism. Typical games are designed such that no matter how players behave, there is always a deterministic outcome. Again, consider the game of rock-paper-scissors as an example. In addition, when assuming that nature is a player, determinism holds, at least on the macro level.
@@ 1.4 Our work
According to which of the three properties we want to keep, there are eight coalition logics. The completeness $\mathsf{CL}$ is proved by Pauly [8], and the completeness of $\mathsf{MCL}$ is shown by Li and Ju [5]. The completeness of the other six logics has not been shown yet.
In this work, we show the completeness of the eight logics by an uniform approach. Generally speaking, our proof approach is as follows. First, we transform every formula to a conjunction of some standard disjunctions. Second, we show a downward validity lemma, which reduces the validity of a standard disjunction to the validity of some formulas with lower modal depth. Third, we show an upward derivability lemma, which reduces the derivability of a standard disjunction to the derivability of some formulas with lower modal depth. Fourth, we show the completeness by induction.
Generally, our approach follows the approach of Goranko and van Drimmelen [3], who showed the completeness of $\mathsf{ATL}$, and Li and Ju [5], although there is much difference between them in detail. Pauly [8] showed the completeness of $\mathsf{CL}$ by the following approach: first, he showed a representation theorem and transformed the semantics for $\mathsf{CL}$ based on concurrent game models to a kind of neighborhood semantics; second, he showed the completeness of $\mathsf{CL}$ with respect to the neighborhood semantics by the Henkin method222 Goranko, Jamroga, and Turrini [4] pointed out that the representation theorem given by Pauly [8] is incorrect, and consequently, the neighborhood semantics given by Pauly [8] is incorrect. They offered the correct representation theorem and neighborhood semantics. However, they showed that the neighborhood semantics given by Pauly [8] determines the same set of valid formulas as the correct neighborhood semantics. . Our approach is different from this approach.
@@
2 Eight coalition logics determined by eight classes of general concurrent game models
@@ 2.1 General concurrent game models and three properties
A general concurrent game frame is a tuple $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$, where:
@@
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame.
@@
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame. We say:
@@
A general concurrent game model is a tuple $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, where:
  • $(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ is a frame;
  • $\mathsf{AP}$ is a countable set of atomic propositions;
@@
For any $\mathsf{X}\in\mathrm{ES}$, a model $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$ is called an $\mathsf{X}$-model if $(\mathtt{ST},\mathtt{AG},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}})$ is an $\mathsf{X}$-frame. Note that $\epsilon$-models are models of $\mathsf{MCL}$ and $\mathsf{S}\mathsf{I}\mathsf{D}$-models are models of $\mathsf{CL}$.
@@ 2.2 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. Define the propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ as usual. Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{A}$ enables $\phi$.
@@ 2.3 Semantics
Let $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$ be a model.
@@
$\mathtt{M},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{M},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{M},s\Vdash\phi$
$\mathtt{M},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{M},s\Vdash\phi$ and $\mathtt{M},s\Vdash\psi$
$\mathtt{M},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is ${\sigma_{\mathrm{C}}}\in\mathtt{aja}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})$, $\mathtt{M},t\Vdash\phi$
@@
$\mathtt{M},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{aja}_{\mathrm{C}}(s)$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{M},t\Vdash\phi$
@@ 2.4 Axiomatic systems
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
@@
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
@@
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
@@
Condition for empty action ($\mathtt{A}\text{-}\mathtt{CEA}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\emptyset]\rangle\top$
Special independence of agents ($\mathtt{A}\text{-}\mathtt{SIA}$): $(\langle[\emptyset]\rangle\phi\land\langle[\mathrm{C}]\rangle\psi)\rightarrow\langle[\mathrm{C}]\rangle(\phi\land\psi)$
Monotonicity ($\mathtt{Mon}$): from $\phi\rightarrow\psi$, we can get $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\psi$,
@@
The formula $\langle[\mathrm{C}]\rangle\top$ defines the class of frames where $\mathrm{C}$ at every state has an available joint action. However, the formula $(\langle[\mathrm{C}]\rangle p\land\langle[\mathrm{D}]\rangle q)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(p\land q)$, where $\mathrm{C}$ and $\mathrm{D}$ are disjoint, does not define the class of frames where at every state, the union of any joint actions of $\mathrm{C}$ and $\mathrm{D}$ is a joint action of $\mathrm{C}\cup\mathrm{D}$; the formula $\langle[\mathrm{C}]\rangle(p\lor q)\rightarrow(\langle[\mathrm{C}]\rangle p\lor\langle[\mathtt{AG}]\rangle q)$ does not define the class of frames where at every state, every available joint action of $\mathtt{AG}$ has a unique outcome state.
In the literature, the formula ¬[]$\neg\langle[\emptyset]\rangle\neg\phi\rightarrow\langle[\mathtt{AG}]\rangle\phi$ is commonly treated as an axiom related to determinism. We want to point out that it is not valid in all deterministic frames, although it is valid in all serial and deterministic frames.

### v2

@@ Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In a recent paper, Li and Ju argued that generally, models of Coalition Logic, concurrent game models, have three too strong assumptions: seriality, independence of agents, and determinism. They presented a Minimal Coalition Logic based on general concurrent game models, which do not have the three assumptions. However, when constructing coalition logics about strategic reasoning in special kinds of situations, we may want to keep some of the assumptions. Thus, studying coalition logics with some of these assumptions makes good sense. In this paper, we show the completeness of these coalition logics in a uniform way.
@@ 1 Introduction
1.1 Coalition Logic and Minimal Coalition Logic
@@
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$, indicating some available joint action of the coalition $\mathrm{A}$ ensures $\phi$. Its models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; at every state, every coalition has some available joint actions; every joint action of a coalition has some possible outcome states. The formula $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{A}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
@@
Recently, Li and Ju [7] argued that, generally, concurrent game models have three too strong assumptions. The first one is seriality: coalitions always have available joint actions. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: joint actions of the grand coalition always have a unique outcome.
Here, we briefly mention some arguments from [7]. Games often terminate when reaching specific states. In terminating states, players do not have available actions intuitively. Thus, seriality might not hold. There are situations where whether a coalition can perform an action is conditional on other agents’ actions at the same time. The independence of agents fails in these situations. Here is an example. There are two agents in a room, $a$ and $b$, and only one chair. Agent $a$ can sit, and agent $b$ can sit, but they cannot sit at the same time. In many situations, some joint actions of all behaving agents have more than one outcome state. The following example, from [12], can illustrate this. A vase stands on a table. There is an agent $a$ who can lift or lower the end of the table. If the table tilts, the vase might fall, and if it falls, it might break. Therefore, determinism might not hold.
Based on general concurrent game models, which do not have the three assumptions, Li and Ju [7] presented a Minimal Coalition Logic $\mathsf{MCL}$.
1.2 Logics between Coalition Logic and Minimal Coalition Logic
Although the three assumptions of concurrent game models are generally too strong, we might want to keep some of them for different reasons when constructing strategic logics.
First, consider seriality. Dropping seriality often makes things complicated. Take Alternating-time Temporal Logic $\mathsf{ATL}$ as an example, where paths, that are infinite sequences of states, play an important role. If we drop seriality, we have to consider finite paths and infinite paths, which is somewhat complicated. Scholars often want to keep seriality by assuming that agents can always perform the special action doing nothing, which seems not harmful.
Second, consider the independence of agents. In many situations, agents’ actions do not interfere with each other, and the independence of agents holds. For example, in the game of rock-paper-scissors, every player can perform three actions: showing rock, paper, and scissors, which do not interfere with each other. As mentioned by Sergot [12], the special actions of attempting to do something do not interfere with each other. Some literature, such as [5] and [6], deals with agency by propositional control: agents are able to control truth values of atomic propositions, and every atomic proposition is controlled by at most one agent. In these settings, the independence of agents holds.
@@ 1.2 Logics between Coalition Logic and Minimal Coalition Logic
Third, consider determinism. Typical games are designed such that no matter how players behave, there is always a deterministic outcome. Again, the game of rock-paper-scissors is an example. In addition, when assuming that nature is a player, determinism holds, at least on the macro level.
According to which of the three properties we want to keep, there are eight coalition logics in total. It makes good sense to study the other six coalition logics.
@@ 1.3 Our work
The completeness $\mathsf{CL}$ is proved by Pauly [10], and the completeness of $\mathsf{MCL}$ is shown by Li and Ju [7]. The completeness of the other six logics has not been shown yet. In this work, we show the completeness of the eight logics in a uniform way by the reduction method.
Generally speaking, to show the completeness of a logic, the reduction method goes through the following four steps. First, we transform a formula to a conjunction of some standard disjunctions. Second, we show a downward validity lemma, which reduces the validity of a standard disjunction to the validity of some formulas with lower modal depth. Third, we show an upward derivability lemma, which reduces the derivability of a standard disjunction to the derivability of some formulas with lower modal depth. Fourth, we show the completeness by induction.
Among the four steps, the first and fourth steps are straightforward, the third step is relatively easy, and the second step, showing the downward validity lemma, is crucial. The key part of the second step is to obtain a pointed model that works.
In showing the completeness of the eight coalition logics, we show the downward validity lemmas in a uniform way, where we use a special kind of abstract structure, called blueprints, to get a pointed model that works.
Pauly [10] showed the completeness of $\mathsf{CL}$ by the following approach: first, he showed a representation theorem and transformed the semantics for $\mathsf{CL}$ based on concurrent game models to some kind of neighborhood semantics; second, he showed the completeness of $\mathsf{CL}$ with respect to the neighborhood semantics by the Henkin method222 Goranko, Jamroga, and Turrini [4] pointed out that the representation theorem given by Pauly [10] is incorrect, and consequently, the neighborhood semantics given by Pauly [10] is incorrect. They offered the correct representation theorem and neighborhood semantics. However, they showed that the neighborhood semantics given by Pauly [10] determines the same set of valid formulas as the correct neighborhood semantics. . Our approach differs from this approach.
Goranko and van Drimmelen [3] showed the completeness of $\mathsf{ATL}$. As mentioned above, $\mathsf{ATL}$ is a temporal extension of $\mathsf{CL}$. Their proof contains a proof for the completeness of $\mathsf{CL}$, which, roughly speaking, is based on (finite) maximal consistent sets. Our proof is different from this proof.
Li and Ju [7] proved the completeness of $\mathsf{MCL}$ by the reduction method, too. The way the downward validity lemmas for the eight coalition logics are shown in this paper is different from how the downward validity lemma for $\mathsf{MCL}$ is shown in [7]. To be more specific, in the latter, we directly constructed a working pointed model (without using blueprints), which heavily depends on the fact that for Minimal Coalition Logic, we do not have to consider seriality, independence of agents, or determinism at all.
@@
2 A family of eight coalition logics
@@ 2.1 General concurrent game models and three properties
A general concurrent game frame is a tuple $\mathfrak{F}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$, where:
@@
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame.
@@
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame. We say:
@@
A general concurrent game model is a tuple $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
  • $(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ is a frame;
@@
For any $\mathsf{X}\in\mathrm{ES}$, a model $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ is called an $\mathsf{X}$-model if $(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ is an $\mathsf{X}$-frame. Note that $\epsilon$-models are models of $\mathsf{MCL}$ and $\mathsf{S}\mathsf{I}\mathsf{D}$-models are models of $\mathsf{CL}$.
@@ 2.2 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$
The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. Define the propositional connectives $\bot,\lor,\rightarrow$ and $\leftrightarrow$ as usual. Define the dual $\hskip-2.5pt\langle\mathrm{C}\rangle\hskip-2.5pt]\phi$ of $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ as $\neg\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\neg\phi$, indicating every available joint action of $\mathrm{A}$ enables $\phi$.
@@ 2.3 Semantics
Let $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a model.
@@
$\mathtt{M},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
$\mathtt{M},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{M},s\Vdash\phi$
$\mathtt{M},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{M},s\Vdash\phi$ and $\mathtt{M},s\Vdash\psi$
$\mathtt{M},s\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is ${\sigma_{\mathrm{C}}}\in\mathtt{aja}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})$, $\mathtt{M},t\Vdash\phi$
@@
$\mathtt{M},s\Vdash[\hskip-2.5pt\langle\mathrm{C}\rangle\hskip-2.5pt]\phi$ $\Leftrightarrow$ for all $\sigma_{\mathrm{C}}\in\mathtt{aja}_{\mathrm{C}}(s)$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{M},t\Vdash\phi$
@@ 2.4 Axiomatic systems
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{C}\cup\mathrm{D}]\hskip-2.0pt\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$
@@
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
@@
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi}$
@@
Condition for empty action ($\mathtt{A}\text{-}\mathtt{CEA}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$
Special independence of agents ($\mathtt{A}\text{-}\mathtt{SIA}$): $(\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle(\phi\land\psi)$
Monotonicity ($\mathtt{Mon}$): from $\phi\rightarrow\psi$, we can get $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\psi$,
@@
The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\top$ defines the class of frames where $\mathrm{C}$ at every state has an available joint action. However, the formula $(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle p\land\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle q)\rightarrow\langle\hskip-2.0pt[\mathrm{C}\cup\mathrm{D}]\hskip-2.0pt\rangle(p\land q)$, where $\mathrm{C}$ and $\mathrm{D}$ are disjoint, does not define the class of frames where at every state, the union of any joint actions of $\mathrm{C}$ and $\mathrm{D}$ is a joint action of $\mathrm{C}\cup\mathrm{D}$; the formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle(p\lor q)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle p\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle q)$ does not define the class of frames where at every state, every available joint action of $\mathtt{AG}$ has a unique outcome state.
In the literature, the formula $\neg\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\phi\rightarrow\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\phi$ is commonly treated as an axiom related to determinism. We want to point out that it is not valid in all deterministic frames, although it is valid in all serial and deterministic frames.